sprockets R Us is selling its bicycle sprockets for 2/5 of a dollar. this price shows a decrease of 1/3 of the original price. what was the original price?
Answer by rfer(16322)

(2/5)/(2/3)
2/5*3/2=6/10=$0.60
.40/2/3=
0.40*3/2=$0.60